OpenStudy (anonymous):

(Newton's Law of Motion) A person stands on a scale in an elevator. When the elevator is at rest, the scale reads 700 N. When the elevator starts to move, the scale reads 600N. (a) Is the elevator going up or down? Show your solution to support your answer. (b) What is its acceleration?

OpenStudy (anonymous):

if elevator going up, acceleration is increase and if it going down ,acceleration is decrease. 700-600=m*a if g=9.8 then a=1.4

OpenStudy (anonymous):

The scale experiences a downwards force of 600N on it from the person, during initial motion; therefore it exerts an equal and opposite force on this person. Since we know the mass of the person (700N/g=70 Kg), let's apply Newton's 2nd law to it: \[-m g + 600N = m a \rightarrow a=(-100N/70Kg)=-1.4 m.s ^{-2}\] The elevator is therefore moving down, as intuitively expected if you've ever been in an elevator. Of course, once the elevator speed becomes constant, we won't be able to tell if we're moving up, down, or even moving!

OpenStudy (anonymous):

Thank you.